Schultz Electronics manufactures two ultra high-definition television models: the Royale which sells for $1,480, and a new model, the Majestic, which sells for $1,320. The production cost computed per unit under traditional costing for each model in 2017 was as follows.

Traditional Costing

Royale

Majestic

Direct materials

$600

$430

Direct labor ($20 per hour)

120

100

Manufacturing overhead ($41 per DLH)

246

205

Total per unit cost

$966

$735


In 2017, Schultz manufactured 25,000 units of the Royale and 10,000 units of the Majestic. The overhead rate of $41 per direct labor hour was determined by dividing total expected manufacturing overhead of $8,131,020 by the total direct labor hours (200,000) for the two models.

Under traditional costing, the gross profit on the models was Royale $514 ($1,480 – $966) and Majestic $585 ($1,320 – $735). Because of this difference, management is considering phasing out the Royale model and increasing the production of the Majestic model.

Before finalizing its decision, management asks Schultz’s controller to prepare an analysis using activity-based costing (ABC). The controller accumulates the following information about overhead for the year ended December 31, 2017.

Activity
Cost Pools

Cost Drivers

Estimated
Overhead

Expected Use of
Cost Drivers

Activity-Based
Overhead Rate

Purchasing Number of orders $1,290,300 39,100 $33/order
Machine setups Number of setups 1,010,820 19,820 $51/setup
Machining Machine hours 5,006,400 119,200 $42/hour
Quality control Number of inspections 823,500 30,500 $27/inspection


The cost drivers used for each product were:

Cost Drivers

Royale

Majestic

Total

Purchase orders 16,900 22,200 39,100
Machine setups 5,820 14,000 19,820
Machine hours 74,600 44,600 119,200
Inspections 11,800 18,700 30,500

(a)

Assign the total 2017 manufacturing overhead costs to the two products using activity-based costing (ABC) and determine the overhead cost per unit. (Round cost per unit to 2 decimal places, e.g. 12.25.)

Royale

Majestic

Total assigned costs

$

$

Cost per unit

$

$
Answer #1
Royale Majestic
Total assigned costs 4306320 3824700
Cost per unit 172.25 382.47
Workings:
Royale Majestic
Purchasing 557700 732600
Machine setups 296820 714000
Machining 3133200 1873200
Quality control 318600 504900
Total assigned costs 4306320 3824700
Divide by units 25000 10000
Cost per unit 172.25 382.47
Royale Majestic
Purchasing =16900*33 =22200*33
Machine setups =5820*51 =14000*51
Machining =74600*42 =44600*42
Quality control =11800*27 =18700*27